# Mathematical Language Models: A Survey

### v3

@@ 7. Conclusions
  • Lu et al. (2023b) Pan Lu, Liang Qiu, Kai-Wei Chang, Ying Nian Wu, Song-Chun Zhu, Tanmay Rajpurohit, Peter Clark, and Ashwin Kalyan. 2023b. Dynamic Prompt Learning via Policy Gradient for Semi-structured Mathematical Reasoning. In The Eleventh International Conference on Learning Representations.
@@
  • Lu et al. (2021b) Pan Lu, Liang Qiu, Jiaqi Chen, Tony Xia, Yizhou Zhao, Wei Zhang, Zhou Yu, Xiaodan Liang, and Song-Chun Zhu. 2021b. IconQA: A New Benchmark for Abstract Diagram Understanding and Visual Language Reasoning. In Thirty-fifth Conference on Neural Information Processing Systems Datasets and Benchmarks Track (Round 2).
@@
  • Lu et al. (2023c) Pan Lu, Liang Qiu, Wenhao Yu, Sean Welleck, and Kai-Wei Chang. 2023c. A Survey of Deep Learning for Mathematical Reasoning. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Anna Rogers, Jordan Boyd-Graber, and Naoaki Okazaki (Eds.). Association for Computational Linguistics, Toronto, Canada, 14605–14631. https://doi.org/10.18653/v1/2023.acl-long.817
@@
  • Luo et al. (2023) Haipeng Luo, Qingfeng Sun, Can Xu, Pu Zhao, Jianguang Lou, Chongyang Tao, Xiubo Geng, Qingwei Lin, Shifeng Chen, and Dongmei Zhang. 2023. Wizardmath: Empowering mathematical reasoning for large language models via reinforced evol-instruct. arXiv preprint arXiv:2308.09583 (2023).
@@
  • Madaan et al. (2023) Aman Madaan, Niket Tandon, Prakhar Gupta, Skyler Hallinan, Luyu Gao, Sarah Wiegreffe, Uri Alon, Nouha Dziri, Shrimai Prabhumoye, Yiming Yang, et al. 2023. Self-refine: Iterative refinement with self-feedback. arXiv preprint arXiv:2303.17651 (2023).
@@
  • Maynez et al. (2020) Joshua Maynez, Shashi Narayan, Bernd Bohnet, and Ryan McDonald. 2020. On Faithfulness and Factuality in Abstractive Summarization. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, Dan Jurafsky, Joyce Chai, Natalie Schluter, and Joel Tetreault (Eds.). Association for Computational Linguistics, Online, 1906–1919. https://doi.org/10.18653/v1/2020.acl-main.173
@@
  • Miao et al. (2023) Ning Miao, Yee Whye Teh, and Tom Rainforth. 2023. Selfcheck: Using llms to zero-shot check their own step-by-step reasoning. arXiv preprint arXiv:2308.00436 (2023).
  • Miao et al. (2020) Shen-Yun Miao, Chao-Chun Liang, and Keh-Yih Su. 2020. A Diverse Corpus for Evaluating and Developing English Math Word Problem Solvers. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics. 975–984.
  • Min et al. (2023) Bonan Min, Hayley Ross, Elior Sulem, Amir Pouran Ben Veyseh, Thien Huu Nguyen, Oscar Sainz, Eneko Agirre, Ilana Heintz, and Dan Roth. 2023. Recent Advances in Natural Language Processing via Large Pre-Trained Language Models: A Survey. ACM Comput. Surv. 56, 2, Article 30 (sep 2023), 40 pages. https://doi.org/10.1145/3605943
  • Min et al. (2022) Sewon Min, Mike Lewis, Luke Zettlemoyer, and Hannaneh Hajishirzi. 2022. MetaICL: Learning to Learn In Context. In Proceedings of the 2022 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies. 2791–2809.
  • Mishra et al. (2022a) Swaroop Mishra, Matthew Finlayson, Pan Lu, Leonard Tang, Sean Welleck, Chitta Baral, Tanmay Rajpurohit, Oyvind Tafjord, Ashish Sabharwal, Peter Clark, et al. 2022a. LILA: A Unified Benchmark for Mathematical Reasoning. In Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing. 5807–5832.
  • Mishra et al. (2022b) Swaroop Mishra, Arindam Mitra, Neeraj Varshney, Bhavdeep Sachdeva, Peter Clark, Chitta Baral, and Ashwin Kalyan. 2022b. NUMGLUE: A Suite of Fundamental yet Challenging Mathematical Reasoning Tasks. In 60th Annual Meeting of the Association for Computational Linguistics, ACL 2022. Association for Computational Linguistics (ACL), 3505–3523.
  • Mitra and Baral (2016) Arindam Mitra and Chitta Baral. 2016. Learning to use formulas to solve simple arithmetic problems. In Proceedings of the 54th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers). 2144–2153.
  • Mo and Xin (2023) Shentong Mo and Miao Xin. 2023. Tree of Uncertain Thoughts Reasoning for Large Language Models. arXiv preprint arXiv:2309.07694 (2023).
@@
  • Naik et al. (2023) Ranjita Naik, Varun Chandrasekaran, Mert Yuksekgonul, Hamid Palangi, and Besmira Nushi. 2023. Diversity of Thought Improves Reasoning Abilities of Large Language Models. arXiv preprint arXiv:2310.07088 (2023).
@@
  • Nelson (1980) Leonard Nelson. 1980. The socratic method. Thinking: The Journal of Philosophy for Children 2, 2 (1980), 34–38.
@@
  • Newell et al. (1957) A. Newell, J. C. Shaw, and H. A. Simon. 1957. Empirical Explorations of the Logic Theory Machine: A Case Study in Heuristic (IRE-AIEE-ACM ’57 (Western)). Association for Computing Machinery, New York, NY, USA, 218–230. https://doi.org/10.1145/1455567.1455605
@@
  • Nogueira et al. (2021) Rodrigo Nogueira, Zhiying Jiang, and Jimmy Lin. 2021. Investigating the limitations of transformers with simple arithmetic tasks. arXiv preprint arXiv:2102.13019 (2021).
@@
  • Noorbakhsh et al. (2021) Kimia Noorbakhsh, Modar Sulaiman, Mahdi Sharifi, Kallol Roy, and Pooyan Jamshidi. 2021. Pretrained Language Models are Symbolic Mathematics Solvers too! arXiv preprint arXiv:2110.03501 (2021).
@@
  • Nye et al. (2021) Maxwell Nye, Anders Johan Andreassen, Guy Gur-Ari, Henryk Michalewski, Jacob Austin, David Bieber, David Dohan, Aitor Lewkowycz, Maarten Bosma, David Luan, Charles Sutton, and Augustus Odena. 2021. Show Your Work: Scratchpads for Intermediate Computation with Language Models. http://arxiv.org/abs/2112.00114 arXiv:2112.00114 [cs].
  • OpenAI (2023) OpenAI. 2023. GPT-4 Technical Report. arXiv:2303.08774 [cs.CL]
@@
  • Paranjape et al. (2023) Bhargavi Paranjape, Scott Lundberg, Sameer Singh, Hannaneh Hajishirzi, Luke Zettlemoyer, and Marco Tulio Ribeiro. 2023. ART: Automatic multi-step reasoning and tool-use for large language models. arXiv:2303.09014 [cs.CL]
@@
  • Parisi et al. (2022) Aaron Parisi, Yao Zhao, and Noah Fiedel. 2022. TALM: Tool Augmented Language Models. arXiv:2205.12255 [cs.CL]
@@
  • Paster et al. (2023) Keiran Paster, Marco Dos Santos, Zhangir Azerbayev, and Jimmy Ba. 2023. OpenWebMath: An Open Dataset of High-Quality Mathematical Web Text. arXiv:2310.06786 [cs.AI]
@@
  • Patel et al. (2021) Arkil Patel, Satwik Bhattamishra, and Navin Goyal. 2021. Are NLP Models really able to Solve Simple Math Word Problems?. In Proceedings of the 2021 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies. 2080–2094.
@@
  • Peng et al. (2021) Shuai Peng, Ke Yuan, Liangcai Gao, and Zhi Tang. 2021. MathBERT: A Pre-Trained Model for Mathematical Formula Understanding. arXiv:2105.00377 [cs]
@@
  • Pitis et al. (2023) Silviu Pitis, Michael R. Zhang, Andrew Wang, and Jimmy Ba. 2023. Boosted Prompt Ensembles for Large Language Models. http://arxiv.org/abs/2304.05970 arXiv:2304.05970 [cs].
@@
  • Polu and Sutskever (2020a) Stanislas Polu and Ilya Sutskever. 2020a. Generative Language Modeling for Automated Theorem Proving. arXiv:2009.03393 [cs.LG]
@@
  • Polu and Sutskever (2020b) Stanislas Polu and Ilya Sutskever. 2020b. Generative Language Modeling for Automated Theorem Proving. CoRR abs/2009.03393 (2020). arXiv:2009.03393 https://arxiv.org/abs/2009.03393
@@
  • Qi et al. (2023) Jingyuan Qi, Zhiyang Xu, Ying Shen, Minqian Liu, dingnan jin, Qifan Wang, and Lifu Huang. 2023. The Art of SOCRATIC QUESTIONING: Recursive Thinking with Large Language Models. In Conference on Empirical Methods in Natural Language Processing. https://api.semanticscholar.org/CorpusID:264935025
@@
  • Qiao et al. (2023) Shuofei Qiao, Yixin Ou, Ningyu Zhang, Xiang Chen, Yunzhi Yao, Shumin Deng, Chuanqi Tan, Fei Huang, and Huajun Chen. 2023. Reasoning with Language Model Prompting: A Survey. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Anna Rogers, Jordan Boyd-Graber, and Naoaki Okazaki (Eds.). Association for Computational Linguistics, Toronto, Canada, 5368–5393. https://doi.org/10.18653/v1/2023.acl-long.294
@@
  • Qin et al. (2020) Jinghui Qin, Lihui Lin, Xiaodan Liang, Rumin Zhang, and Liang Lin. 2020. Semantically-Aligned Universal Tree-Structured Solver for Math Word Problems. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP). 3780–3789.
@@
  • Rawte et al. (2023) Vipula Rawte, Amit Sheth, and Amitava Das. 2023. A survey of hallucination in large foundation models. arXiv preprint arXiv:2309.05922 (2023).
@@
  • Roy and Roth (2015) Subhro Roy and Dan Roth. 2015. Solving General Arithmetic Word Problems. In Proceedings of the 2015 Conference on Empirical Methods in Natural Language Processing. 1743–1752.
@@
  • Roy and Roth (2017) Subhro Roy and Dan Roth. 2017. Unit dependency graph and its application to arithmetic word problem solving. In Proceedings of the AAAI conference on artificial intelligence, Vol. 31.
@@
  • Roy and Roth (2018) Subhro Roy and Dan Roth. 2018. Mapping to declarative knowledge for word problem solving. Transactions of the Association for Computational Linguistics 6 (2018), 159–172.
@@
  • Roy et al. (2015) Subhro Roy, Tim Vieira, and Dan Roth. 2015. Reasoning about quantities in natural language. Transactions of the Association for Computational Linguistics 3 (2015), 1–13.
@@
  • Rudnicki (1992) Piotr Rudnicki. 1992. An overview of the Mizar project. In Proceedings of the 1992 Workshop on Types for Proofs and Programs. Citeseer, 311–330.
@@
  • Saxton et al. (2019) David Saxton, Edward Grefenstette, Felix Hill, and Pushmeet Kohli. 2019. Analysing Mathematical Reasoning Abilities of Neural Models. In International Conference on Learning Representations.
@@
  • Scharpf et al. (2022) Philipp Scharpf, Moritz Schubotz, and Bela Gipp. 2022. Mining mathematical documents for question answering via unsupervised formula labeling. In Proceedings of the 22nd ACM/IEEE Joint Conference on Digital Libraries. 1–11.
@@
  • Schick et al. (2023) Timo Schick, Jane Dwivedi-Yu, Roberto Dessì, Roberta Raileanu, Maria Lomeli, Luke Zettlemoyer, Nicola Cancedda, and Thomas Scialom. 2023. Toolformer: Language Models Can Teach Themselves to Use Tools. arXiv:2302.04761 [cs.CL]
@@
  • Schulman et al. (2017) John Schulman, Filip Wolski, Prafulla Dhariwal, Alec Radford, and Oleg Klimov. 2017. Proximal policy optimization algorithms. arXiv preprint arXiv:1707.06347 (2017).
@@
  • Sel et al. (2023) Bilgehan Sel, Ahmad Al-Tawaha, Vanshaj Khattar, Lu Wang, Ruoxi Jia, and Ming Jin. 2023. Algorithm of thoughts: Enhancing exploration of ideas in large language models. arXiv preprint arXiv:2308.10379 (2023).
@@
  • Seo et al. (2015) Minjoon Seo, Hannaneh Hajishirzi, Ali Farhadi, Oren Etzioni, and Clint Malcolm. 2015. Solving Geometry Problems: Combining Text and Diagram Interpretation. In Proceedings of the 2015 Conference on Empirical Methods in Natural Language Processing, Lluís Màrquez, Chris Callison-Burch, and Jian Su (Eds.). Association for Computational Linguistics, Lisbon, Portugal, 1466–1476. https://doi.org/10.18653/v1/D15-1171
@@
  • Shen et al. (2021) Jianhao Shen, Yichun Yin, Lin Li, Lifeng Shang, Xin Jiang, Ming Zhang, and Qun Liu. 2021. Generate & Rank: A Multi-task Framework for Math Word Problems. http://arxiv.org/abs/2109.03034 arXiv:2109.03034 [cs].
@@
  • Shi et al. (2022) Freda Shi, Mirac Suzgun, Markus Freitag, Xuezhi Wang, Suraj Srivats, Soroush Vosoughi, Hyung Won Chung, Yi Tay, Sebastian Ruder, Denny Zhou, et al. 2022. Language models are multilingual chain-of-thought reasoners. In The Eleventh International Conference on Learning Representations.
@@
  • Shi et al. (2015) Shuming Shi, Yuehui Wang, Chin-Yew Lin, Xiaojiang Liu, and Yong Rui. 2015. Automatically solving number word problems by semantic parsing and reasoning. In Proceedings of the 2015 conference on empirical methods in natural language processing. 1132–1142.
@@
  • Shinn et al. (2023) Noah Shinn, Federico Cassano, Ashwin Gopinath, Karthik R Narasimhan, and Shunyu Yao. 2023. Reflexion: Language agents with verbal reinforcement learning. In Thirty-seventh Conference on Neural Information Processing Systems.
@@
  • Shridhar et al. (2023) Kumar Shridhar, Harsh Jhamtani, Hao Fang, Benjamin Van Durme, Jason Eisner, and Patrick Xia. 2023. SCREWS: A Modular Framework for Reasoning with Revisions. arXiv preprint arXiv:2309.13075 (2023).
@@
  • Shridhar et al. (2022) Kumar Shridhar, Jakub Macina, Mennatallah El-Assady, Tanmay Sinha, Manu Kapur, and Mrinmaya Sachan. 2022. Automatic Generation of Socratic Subquestions for Teaching Math Word Problems. arXiv preprint arXiv:2211.12835 (2022).
@@
  • Shum et al. (2023) KaShun Shum, Shizhe Diao, and Tong Zhang. 2023. Automatic Prompt Augmentation and Selection with Chain-of-Thought from Labeled Data. http://arxiv.org/abs/2302.12822 arXiv:2302.12822 [cs].
  • Slind and Norrish (2008) Konrad Slind and Michael Norrish. 2008. A brief overview of HOL4. In International Conference on Theorem Proving in Higher Order Logics. Springer, 28–32.
@@
  • Spithourakis et al. (2016) Georgios Spithourakis, Isabelle Augenstein, and Sebastian Riedel. 2016. Numerically Grounded Language Models for Semantic Error Correction. In Proceedings of the 2016 Conference on Empirical Methods in Natural Language Processing. 987–992.
@@
  • Spithourakis and Riedel (2018) GP Spithourakis and S Riedel. 2018. Numeracy for language models: Evaluating and improving their ability to predict numbers. In ACL 2018-56th Annual Meeting of the Association for Computational Linguistics, Proceedings of the Conference (Long Papers), Vol. 56. Association for Computational Linguistics, 2104–2115.
@@
  • Sun et al. (2023) Haotian Sun, Yuchen Zhuang, Lingkai Kong, Bo Dai, and Chao Zhang. 2023. AdaPlanner: Adaptive Planning from Feedback with Language Models. arXiv preprint arXiv:2305.16653 (2023).

### v4

@@ Acknowledgements.
  • Lu et al. (2023c) Pan Lu, Liang Qiu, Kai-Wei Chang, Ying Nian Wu, Song-Chun Zhu, Tanmay Rajpurohit, Peter Clark, and Ashwin Kalyan. 2023c. Dynamic Prompt Learning via Policy Gradient for Semi-structured Mathematical Reasoning. In ICLR.
@@
  • Lu et al. (2021b) Pan Lu, Liang Qiu, Jiaqi Chen, Tony Xia, Yizhou Zhao, Wei Zhang, Zhou Yu, Xiaodan Liang, and Song-Chun Zhu. 2021b. IconQA: A New Benchmark for Abstract Diagram Understanding and Visual Language Reasoning. In NeurIPS Datasets and Benchmarks Track.
@@
  • Lu et al. (2023d) Pan Lu, Liang Qiu, Wenhao Yu, Sean Welleck, and Kai-Wei Chang. 2023d. A Survey of Deep Learning for Mathematical Reasoning. In ACL, Anna Rogers, Jordan Boyd-Graber, and Naoaki Okazaki (Eds.). 14605–14631.
@@
  • Luo et al. (2023) Haipeng Luo, Qingfeng Sun, Can Xu, Pu Zhao, Jianguang Lou, Chongyang Tao, Xiubo Geng, Qingwei Lin, Shifeng Chen, and Dongmei Zhang. 2023. Wizardmath: Empowering mathematical reasoning for large language models via reinforced evol-instruct. arXiv (2023).
  • Luo et al. (2024) Liangchen Luo, Yinxiao Liu, Rosanne Liu, Samrat Phatale, Harsh Lara, Yunxuan Li, Lei Shu, Yun Zhu, Lei Meng, Jiao Sun, et al. 2024. Improve Mathematical Reasoning in Language Models by Automated Process Supervision. arXiv (2024).
  • Luong et al. (2024) Trung Quoc Luong, Xinbo Zhang, Zhanming Jie, Peng Sun, Xiaoran Jin, and Hang Li. 2024. Reft: Reasoning with reinforced fine-tuning. arXiv (2024).
  • Ma et al. (2023) Qianli Ma, Haotian Zhou, Tingkai Liu, Jianbo Yuan, Pengfei Liu, Yang You, and Hongxia Yang. 2023. Let’s reward step by step: Step-Level reward model as the Navigators for Reasoning. arXiv (2023).
  • Macina et al. (2023) Jakub Macina, Nico Daheim, Sankalan Pal Chowdhury, Tanmay Sinha, Manu Kapur, Iryna Gurevych, and Mrinmaya Sachan. 2023. Mathdial: A dialogue tutoring dataset with rich pedagogical properties grounded in math reasoning problems. arXiv (2023).
@@
  • Madaan et al. (2023) Aman Madaan, Niket Tandon, Prakhar Gupta, Skyler Hallinan, Luyu Gao, Sarah Wiegreffe, Uri Alon, Nouha Dziri, Shrimai Prabhumoye, Yiming Yang, et al. 2023. Self-refine: Iterative refinement with self-feedback. arXiv (2023).
@@
  • Maynez et al. (2020) Joshua Maynez, Shashi Narayan, Bernd Bohnet, and Ryan McDonald. 2020. On Faithfulness and Factuality in Abstractive Summarization. In ACL. 1906–1919.
@@
  • Miao et al. (2023) Ning Miao, Yee Whye Teh, and Tom Rainforth. 2023. Selfcheck: Using llms to zero-shot check their own step-by-step reasoning. arXiv (2023).
  • Miao et al. (2020) Shen-Yun Miao, Chao-Chun Liang, and Keh-Yih Su. 2020. A Diverse Corpus for Evaluating and Developing English Math Word Problem Solvers. In ACL. 975–984.
  • Min et al. (2023) Bonan Min, Hayley Ross, Elior Sulem, Amir Pouran Ben Veyseh, Thien Huu Nguyen, Oscar Sainz, Eneko Agirre, Ilana Heintz, and Dan Roth. 2023. Recent Advances in Natural Language Processing via Large Pre-Trained Language Models: A Survey. ACM Comput. Surv. 56, 2 (2023).
  • Min et al. (2022) Sewon Min, Mike Lewis, Luke Zettlemoyer, and Hannaneh Hajishirzi. 2022. MetaICL: Learning to Learn In Context. In NAACL. 2791–2809.
  • Mishra et al. (2022a) Swaroop Mishra, Matthew Finlayson, Pan Lu, Leonard Tang, Sean Welleck, Chitta Baral, Tanmay Rajpurohit, Oyvind Tafjord, Ashish Sabharwal, Peter Clark, et al. 2022a. LILA: A Unified Benchmark for Mathematical Reasoning. In EMNLP. 5807–5832.
  • Mishra et al. (2022b) Swaroop Mishra, Arindam Mitra, Neeraj Varshney, Bhavdeep Sachdeva, Peter Clark, Chitta Baral, and Ashwin Kalyan. 2022b. NUMGLUE: A Suite of Fundamental yet Challenging Mathematical Reasoning Tasks. In ACL. 3505–3523.
  • Mitra and Baral (2016) Arindam Mitra and Chitta Baral. 2016. Learning to use formulas to solve simple arithmetic problems. In ACL. 2144–2153.
  • Mo and Xin (2023) Shentong Mo and Miao Xin. 2023. Tree of Uncertain Thoughts Reasoning for Large Language Models. arXiv (2023).
@@
  • Naik et al. (2023) Ranjita Naik, Varun Chandrasekaran, Mert Yuksekgonul, Hamid Palangi, and Besmira Nushi. 2023. Diversity of Thought Improves Reasoning Abilities of Large Language Models. arXiv (2023).
@@
  • Nelson (1980) Leonard Nelson. 1980. The socratic method. Thinking: The Journal of Philosophy for Children 2, 2, 34–38.
@@
  • Newell et al. (1957) A. Newell, J. C. Shaw, and H. A. Simon. 1957. Empirical Explorations of the Logic Theory Machine: A Case Study in Heuristic. 218–230.
@@
  • Nogueira et al. (2021) Rodrigo Nogueira, Zhiying Jiang, and Jimmy Lin. 2021. Investigating the limitations of transformers with simple arithmetic tasks. arXiv (2021).
@@
  • Noorbakhsh et al. (2021) Kimia Noorbakhsh, Modar Sulaiman, Mahdi Sharifi, Kallol Roy, and Pooyan Jamshidi. 2021. Pretrained Language Models are Symbolic Mathematics Solvers too! arXiv (2021).
@@
  • Nye et al. (2021) Maxwell Nye, Anders Johan Andreassen, Guy Gur-Ari, Henryk Michalewski, Jacob Austin, David Bieber, David Dohan, Aitor Lewkowycz, Maarten Bosma, David Luan, Charles Sutton, and Augustus Odena. 2021. Show Your Work: Scratchpads for Intermediate Computation with Language Models.
  • Olausson et al. (2023) Theo X Olausson, Alex Gu, Benjamin Lipkin, Cedegao E Zhang, Armando Solar-Lezama, Joshua B Tenenbaum, and Roger Levy. 2023. LINC: A neurosymbolic approach for logical reasoning by combining language models with first-order logic provers. arXiv (2023).
  • OpenAI (2023) OpenAI. 2023. GPT-4 Technical Report. arXiv:2303.08774
  • OpenAI (2024) OpenAI. 2024. OpenAI O1 System Card. https://openai.com/index/openai-o1-system-card/
@@
  • Paranjape et al. (2023) Bhargavi Paranjape, Scott Lundberg, Sameer Singh, Hannaneh Hajishirzi, Luke Zettlemoyer, and Marco Tulio Ribeiro. 2023. ART: Automatic multi-step reasoning and tool-use for large language models. arXiv:2303.09014
@@
  • Parisi et al. (2022) Aaron Parisi, Yao Zhao, and Noah Fiedel. 2022. TALM: Tool Augmented Language Models. arXiv:2205.12255
@@
  • Paster et al. (2023) Keiran Paster, Marco Dos Santos, Zhangir Azerbayev, and Jimmy Ba. 2023. OpenWebMath: An Open Dataset of High-Quality Mathematical Web Text. arXiv:2310.06786
@@
  • Patel et al. (2021) Arkil Patel, Satwik Bhattamishra, and Navin Goyal. 2021. Are NLP Models really able to Solve Simple Math Word Problems?. In NAACL. 2080–2094.
@@
  • Peng et al. (2021) Shuai Peng, Ke Yuan, Liangcai Gao, and Zhi Tang. 2021. MathBERT: A Pre-Trained Model for Mathematical Formula Understanding.
@@
  • Pitis et al. (2023) Silviu Pitis, Michael R Zhang, Andrew Wang, and Jimmy Ba. 2023. Boosted prompt ensembles for large language models. arXiv (2023).
@@
  • Polu and Sutskever (2020a) Stanislas Polu and Ilya Sutskever. 2020a. Generative Language Modeling for Automated Theorem Proving. arXiv:2009.03393
@@
  • Polu and Sutskever (2020b) Stanislas Polu and Ilya Sutskever. 2020b. Generative Language Modeling for Automated Theorem Proving. CoRR abs/2009.03393 (2020).
@@
  • Qi et al. (2023) Jingyuan Qi, Zhiyang Xu, Ying Shen, Minqian Liu, dingnan jin, Qifan Wang, and Lifu Huang. 2023. The Art of SOCRATIC QUESTIONING: Recursive Thinking with Large Language Models. In EMNLP.
  • Qian et al. (2023) Cheng Qian, Chi Han, Yi R Fung, Yujia Qin, Zhiyuan Liu, and Heng Ji. 2023. Creator: Tool creation for disentangling abstract and concrete reasoning of large language models. arXiv (2023).
  • Qiao et al. (2024) Runqi Qiao, Qiuna Tan, Guanting Dong, Minhui Wu, Chong Sun, Xiaoshuai Song, Zhuoma GongQue, Shanglin Lei, Zhe Wei, Miaoxuan Zhang, Runfeng Qiao, Yifan Zhang, Xiao Zong, Yida Xu, Muxi Diao, Zhimin Bao, Chen Li, and Honggang Zhang. 2024. We-Math: Does Your Large Multimodal Model Achieve Human-like Mathematical Reasoning? arXiv:2407.01284
@@
  • Qiao et al. (2023) Shuofei Qiao, Yixin Ou, Ningyu Zhang, Xiang Chen, Yunzhi Yao, Shumin Deng, Chuanqi Tan, Fei Huang, and Huajun Chen. 2023. Reasoning with Language Model Prompting: A Survey. In ACL. 5368–5393.
@@
  • Qin et al. (2020) Jinghui Qin, Lihui Lin, Xiaodan Liang, Rumin Zhang, and Liang Lin. 2020. Semantically-Aligned Universal Tree-Structured Solver for Math Word Problems. In EMNLP. 3780–3789.
@@
  • Rafailov et al. (2024) Rafael Rafailov, Archit Sharma, Eric Mitchell, Christopher D Manning, Stefano Ermon, and Chelsea Finn. 2024. Direct preference optimization: Your language model is secretly a reward model. NeurIPS 36 (2024).
  • Rawte et al. (2023) Vipula Rawte, Amit Sheth, and Amitava Das. 2023. A survey of hallucination in large foundation models. arXiv (2023).
@@
  • Roy and Roth (2015) Subhro Roy and Dan Roth. 2015. Solving General Arithmetic Word Problems. In EMNLP. 1743–1752.
@@
  • Roy and Roth (2017) Subhro Roy and Dan Roth. 2017. Unit dependency graph and its application to arithmetic word problem solving. In AAAI, Vol. 31.
@@
  • Roy and Roth (2018) Subhro Roy and Dan Roth. 2018. Mapping to declarative knowledge for word problem solving. TACL 6 (2018), 159–172.
@@
  • Roy et al. (2015) Subhro Roy, Tim Vieira, and Dan Roth. 2015. Reasoning about quantities in natural language. TACL 3 (2015), 1–13.
@@
  • Rudnicki (1992) Piotr Rudnicki. 1992. An overview of the Mizar project. In Proceedings of the 1992 Workshop on Types for Proofs and Programs. 311–330.
@@
  • Saxton et al. (2019) David Saxton, Edward Grefenstette, Felix Hill, and Pushmeet Kohli. 2019. Analysing Mathematical Reasoning Abilities of Neural Models. In ICLR.
@@
  • Scharpf et al. (2022) Philipp Scharpf, Moritz Schubotz, and Bela Gipp. 2022. Mining mathematical documents for question answering via unsupervised formula labeling. In ICDL. 1–11.
@@
  • Schick et al. (2023) Timo Schick, Jane Dwivedi-Yu, Roberto Dessì, Roberta Raileanu, Maria Lomeli, Luke Zettlemoyer, Nicola Cancedda, and Thomas Scialom. 2023. Toolformer: Language Models Can Teach Themselves to Use Tools. arXiv:2302.04761
@@
  • Schulman et al. (2017) John Schulman, Filip Wolski, Prafulla Dhariwal, Alec Radford, and Oleg Klimov. 2017. Proximal policy optimization algorithms. arXiv (2017).
@@
  • Sel et al. (2023) Bilgehan Sel, Ahmad Al-Tawaha, Vanshaj Khattar, Lu Wang, Ruoxi Jia, and Ming Jin. 2023. Algorithm of thoughts: Enhancing exploration of ideas in large language models. arXiv (2023).
@@
  • Seo et al. (2015) Minjoon Seo, Hannaneh Hajishirzi, Ali Farhadi, Oren Etzioni, and Clint Malcolm. 2015. Solving Geometry Problems: Combining Text and Diagram Interpretation. In EMNLP. 1466–1476.
@@
  • Shen et al. (2021) Jianhao Shen, Yichun Yin, Lin Li, Lifeng Shang, Xin Jiang, Ming Zhang, and Qun Liu. 2021. Generate & Rank: A Multi-task Framework for Math Word Problems.
@@
  • Shi et al. (2022) Freda Shi, Mirac Suzgun, Markus Freitag, Xuezhi Wang, Suraj Srivats, Soroush Vosoughi, Hyung Won Chung, Yi Tay, Sebastian Ruder, Denny Zhou, et al. 2022. Language models are multilingual chain-of-thought reasoners. In ICLR.
@@
  • Shi et al. (2015) Shuming Shi, Yuehui Wang, Chin-Yew Lin, Xiaojiang Liu, and Yong Rui. 2015. Automatically solving number word problems by semantic parsing and reasoning. In EMNLP. 1132–1142.
  • Shi et al. (2024) Wenhao Shi, Zhiqiang Hu, Yi Bin, Junhua Liu, Yang Yang, See-Kiong Ng, Lidong Bing, and Roy Ka-Wei Lee. 2024. Math-LLaVA: Bootstrapping Mathematical Reasoning for Multimodal Large Language Models.
@@
  • Shinn et al. (2023) Noah Shinn, Federico Cassano, Ashwin Gopinath, Karthik R Narasimhan, and Shunyu Yao. 2023. Reflexion: Language agents with verbal reinforcement learning. In NeurIPS.
@@
  • Shridhar et al. (2023) Kumar Shridhar, Harsh Jhamtani, Hao Fang, Benjamin Van Durme, Jason Eisner, and Patrick Xia. 2023. SCREWS: A Modular Framework for Reasoning with Revisions. arXiv (2023).
@@
  • Shridhar et al. (2022) Kumar Shridhar, Jakub Macina, Mennatallah El-Assady, Tanmay Sinha, Manu Kapur, and Mrinmaya Sachan. 2022. Automatic Generation of Socratic Subquestions for Teaching Math Word Problems. arXiv (2022).
@@
  • Shum et al. (2023) Kashun Shum, Shizhe Diao, and Tong Zhang. 2023. Automatic Prompt Augmentation and Selection with Chain-of-Thought from Labeled Data. In EMNLP. 12113–12139.
  • Slind and Norrish (2008) Konrad Slind and Michael Norrish. 2008. A brief overview of HOL4. In TPHOLs. 28–32.
  • Song et al. (2024) Peiyang Song, Kaiyu Yang, and Anima Anandkumar. 2024. Towards large language models as copilots for theorem proving in lean. arXiv (2024).
@@
  • Spithourakis et al. (2016) Georgios Spithourakis, Isabelle Augenstein, and Sebastian Riedel. 2016. Numerically Grounded Language Models for Semantic Error Correction. In EMNLP. 987–992.
@@
  • Spithourakis and Riedel (2018) GP Spithourakis and S Riedel. 2018. Numeracy for language models: Evaluating and improving their ability to predict numbers. In ACL, Vol. 56. 2104–2115.
@@
  • Sun et al. (2023) Haotian Sun, Yuchen Zhuang, Lingkai Kong, Bo Dai, and Chao Zhang. 2023. AdaPlanner: Adaptive Planning from Feedback with Language Models. arXiv (2023).